OpenStudy (andijo76):

listed below are the budgets ( in millions of dollars) and the gross receipts ( in millions of dollars for the randomly selected movies construct a scatterplot find the value. of the linear correlation coefficient r and find the P value using a=0.05 is there sufficient evidence to conclude that there is a linear correlation between budgets and gross receipts? do the results change if the actual budgets listed are $63,000.000 $87,000.000 $49,000.000 and so on i have the scatterplot figured out i will show what i have so far

OpenStudy (andijo76):

budget(x) 63,87,49,31,193,96,91 gross(y) 61,68,45,62,536,142,48 budget (x) =63,87,49,31,193,96,91=610 gross(y)= 61,68,45,62,536,142,48=962 x*y= 3843,5916,2205,1922,103448,13632,4368=130966 x^2= 3962,7569,2401,961,37249,9216,8281=70608 y^2= 3721,4624,2025,3844,287296,20164,2304=323978

OpenStudy (andijo76):

x=Ex/n=610/7=87y=Ey/n=962/7=137

OpenStudy (andijo76):



OpenStudy (andijo76):

this is far as i got

OpenStudy (andijo76):

i dont know if it is correct